\begin{thebibliography}{}

\bibitem[\protect\citename{Aitchison \& Lauder, }1985]{doi:10.2307/2347365}
Aitchison, J., \& Lauder, I.~J. 1985.
\newblock Kernel Density Estimation for Compositional Data.
\newblock {\em Journal of the Royal Statistical Society. Series C (Applied
  Statistics)}, {\bf 34}(2), 129--137.
\newblock \href{https://doi.org/10.2307/2347365}{doi:10.2307/2347365}.

\bibitem[\protect\citename{Arenbaev, }1976]{MR0478288}
Arenbaev, N.~K. 1976.
\newblock Asymptotic behavior of the multinomial distribution.
\newblock {\em Teor. Veroyatnost. i Primenen.}, {\bf 21}(4), 826--831.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR0478288}{MR0478288}.

\bibitem[\protect\citename{Babu {\em et~al.}, }2002]{MR1910059}
Babu, G.~J., Canty, A.~J., \& Chaubey, Y.~P. 2002.
\newblock Application of {B}ernstein polynomials for smooth estimation of a
  distribution and density function.
\newblock {\em J. Statist. Plann. Inference}, {\bf 105}(2), 377--392.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR1910059}{MR1910059}.

\bibitem[\protect\citename{Belalia {\em et~al.}, }2017]{MR3630225}
Belalia, M., Bouezmarni, T., \& Leblanc, A. 2017.
\newblock Smooth conditional distribution estimators using {B}ernstein
  polynomials.
\newblock {\em Comput. Statist. Data Anal.}, {\bf 111}, 166--182.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR3630225}{MR3630225}.

\bibitem[\protect\citename{Belalia {\em et~al.}, }2019]{MR3983257}
Belalia, M., Bouezmarni, T., \& Leblanc, A. 2019.
\newblock Bernstein conditional density estimation with application to
  conditional distribution and regression functions.
\newblock {\em J. Korean Statist. Soc.}, {\bf 48}(3), 356--383.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR3983257}{MR3983257}.

\bibitem[\protect\citename{Bouezmarni \& Rolin, }2003]{MR1985506}
Bouezmarni, T., \& Rolin, J.-M. 2003.
\newblock Consistency of the beta kernel density function estimator.
\newblock {\em Canad. J. Statist.}, {\bf 31}(1), 89--98.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR1985506}{MR1985506}.

\bibitem[\protect\citename{Bouezmarni \& Rolin, }2007]{MR2351744}
Bouezmarni, T., \& Rolin, J.-M. 2007.
\newblock Bernstein estimator for unbounded density function.
\newblock {\em J. Nonparametr. Stat.}, {\bf 19}(3), 145--161.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR2351744}{MR2351744}.

\bibitem[\protect\citename{Bouezmarni {\em et~al.}, }2007]{MR2395599}
Bouezmarni, T., M., Mesfioui, \& Rolin, J.~M. 2007.
\newblock $L_1$-rate of convergence of smoothed histogram.
\newblock {\em Statist. Probab. Lett.}, {\bf 77}(14), 1497--1504.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR2395599}{MR2395599}.

\bibitem[\protect\citename{Brown \& Chen, }1999]{MR1685301}
Brown, B.~M., \& Chen, S.~X. 1999.
\newblock Beta-{B}ernstein smoothing for regression curves with compact
  support.
\newblock {\em Scand. J. Statist.}, {\bf 26}(1), 47--59.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR1685301}{MR1685301}.

\bibitem[\protect\citename{Chen, }1999]{MR1718494}
Chen, S.~X. 1999.
\newblock Beta kernel estimators for density functions.
\newblock {\em Comput. Statist. Data Anal.}, {\bf 31}(2), 131--145.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR1718494}{MR1718494}.

\bibitem[\protect\citename{Chen, }2000]{MR1742101}
Chen, S.~X. 2000.
\newblock Beta kernel smoothers for regression curves.
\newblock {\em Statist. Sinica}, {\bf 10}(1), 73--91.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR1742101}{MR1742101}.

\bibitem[\protect\citename{Curtis \& Ghosh, }2011]{MR2782409}
Curtis, S.~M., \& Ghosh, S.~K. 2011.
\newblock A variable selection approach to monotonic regression with
  {B}ernstein polynomials.
\newblock {\em J. Appl. Stat.}, {\bf 38}(5), 961--976.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR2782409}{MR2782409}.

\bibitem[\protect\citename{Gawronski, }1985]{MR0791719}
Gawronski, W. 1985.
\newblock Strong laws for density estimators of {B}ernstein type.
\newblock {\em Period. Math. Hungar}, {\bf 16}(1), 23--43.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR0791719}{MR0791719}.

\bibitem[\protect\citename{Gawronski \& Stadtm\"uller, }1981]{MR0638651}
Gawronski, W., \& Stadtm\"uller, U. 1981.
\newblock Smoothing histograms by means of lattice and continuous
  distributions.
\newblock {\em Metrika}, {\bf 28}(3), 155--164.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR0638651}{MR0638651}.

\bibitem[\protect\citename{Ghosal, }2001]{MR1873330}
Ghosal, S. 2001.
\newblock Convergence rates for density estimation with {B}ernstein
  polynomials.
\newblock {\em Ann. Statist.}, {\bf 29}(5), 1264--1280.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR1873330}{MR1873330}.

\bibitem[\protect\citename{Guan, }2016]{MR3488598}
Guan, Z. 2016.
\newblock Efficient and robust density estimation using {B}ernstein type
  polynomials.
\newblock {\em J. Nonparametr. Stat.}, {\bf 28}(2), 250--271.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR3488598}{MR3488598}.

\bibitem[\protect\citename{Guan, }2017]{MR3740722}
Guan, Z. 2017.
\newblock Bernstein polynomial model for grouped continuous data.
\newblock {\em J. Nonparametr. Stat.}, {\bf 29}(4), 831--848.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR3740722}{MR3740722}.

\bibitem[\protect\citename{Igarashi \& Kakizawa, }2014]{MR3174309}
Igarashi, G., \& Kakizawa, Y. 2014.
\newblock On improving convergence rate of {B}ernstein polynomial density
  estimator.
\newblock {\em J. Nonparametr. Stat.}, {\bf 26}(1), 61--84.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR3174309}{MR3174309}.

\bibitem[\protect\citename{Kakizawa, }2004]{MR2068610}
Kakizawa, Y. 2004.
\newblock Bernstein polynomial probability density estimation.
\newblock {\em J. Nonparametr. Stat.}, {\bf 16}(5), 709--729.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR2068610}{MR2068610}.

\bibitem[\protect\citename{Leblanc, }2009]{MR2488150}
Leblanc, A. 2009.
\newblock Chung-{S}mirnov property for {B}ernstein estimators of distribution
  functions.
\newblock {\em J. Nonparametr. Stat.}, {\bf 21}(2), 133--142.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR2488150}{MR2488150}.

\bibitem[\protect\citename{Leblanc, }2010]{MR2662607}
Leblanc, A. 2010.
\newblock A bias-reduced approach to density estimation using {B}ernstein
  polynomials.
\newblock {\em J. Nonparametr. Stat.}, {\bf 22}(3-4), 459--475.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR2662607}{MR2662607}.

\bibitem[\protect\citename{Leblanc, }2012a]{MR2960952}
Leblanc, A. 2012a.
\newblock On estimating distribution functions using {B}ernstein polynomials.
\newblock {\em Ann. Inst. Statist. Math.}, {\bf 64}(5), 919--943.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR2960952}{MR2960952}.

\bibitem[\protect\citename{Leblanc, }2012b]{MR2925964}
Leblanc, A. 2012b.
\newblock On the boundary properties of {B}ernstein polynomial estimators of
  density and distribution functions.
\newblock {\em J. Statist. Plann. Inference}, {\bf 142}(10), 2762--2778.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR2925964}{MR2925964}.

\bibitem[\protect\citename{Lu, }2015]{MR3412755}
Lu, L. 2015.
\newblock On the uniform consistency of the {B}ernstein density estimator.
\newblock {\em Statist. Probab. Lett.}, {\bf 107}, 52--61.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR3412755}{MR3412755}.

\bibitem[\protect\citename{Mosimann, }1962]{MR143299}
Mosimann, J.~E. 1962.
\newblock On the compound multinomial distribution, the multivariate {$\beta
  $}-distribution, and correlations among proportions.
\newblock {\em Biometrika}, {\bf 49}, 65--82.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR143299}{MR143299}.

\bibitem[\protect\citename{Newcomer, }2008]{Newcomer2008phd}
Newcomer, J.~T. 2008.
\newblock {\em Estimation procedures for multinomial models with
  overdispersion}.
\newblock Ph{D} thesis, University of Maryland.

\bibitem[\protect\citename{Newcomer {\em et~al.},
  }2008]{Newcomer_et_al_2008_tech_report}
Newcomer, J.~T., Neerchal, N.~K., \& Morel, J.~G. 2008.
\newblock Computation of higher order moments from two multinomial
  overdispersion likelihood models.
\newblock {\em Preprint},  1--11.
\newblock
  \url{http://www.math.umbc.edu/~kogan/technical_papers/2008/Newcomer_Nagaraj_Morel.pdf}.

\bibitem[\protect\citename{Ouimet, }2018]{MR3825458}
Ouimet, F. 2018.
\newblock Complete monotonicity of multinomial probabilities and its
  application to {B}ernstein estimators on the simplex.
\newblock {\em J. Math. Anal. Appl.}, {\bf 466}(2), 1609--1617.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR3825458}{MR3825458}.

\bibitem[\protect\citename{Ouimet, }2019]{Ouimet2019phd}
Ouimet, F. 2019.
\newblock {\em Extremes of log-correlated random fields and the {R}iemann-zeta
  function, and some asymptotic results for various estimators in statistics}.
\newblock Ph{D} thesis, Universit\'e de Montr\'eal.
\newblock \url{http://hdl.handle.net/1866/22667}.

\bibitem[\protect\citename{Ouimet, }2020a]{arXiv:2002.07758}
Ouimet, F. 2020a.
\newblock Asymptotic properties of {B}ernstein estimators on the simplex.
\newblock {\em Preprint},  1--27.
\newblock \href{https://arxiv.org/abs/2002.07758}{arXiv:2002.07758}.

\bibitem[\protect\citename{Ouimet, }2020b]{arXiv:2006.11756}
Ouimet, F. 2020b.
\newblock Asymptotic properties of {B}ernstein estimators on the simplex. Part
  2: the boundary case.
\newblock {\em Preprint},  1--23.
\newblock \href{https://arxiv.org/abs/2006.11756}{arXiv:2006.11756}.

\bibitem[\protect\citename{Ouimet, }2020c]{arXiv:2002.06956}
Ouimet, F. 2020c.
\newblock Density estimation using Dirichlet kernels.
\newblock {\em Preprint},  1--39.
\newblock \href{https://arxiv.org/abs/2002.06956}{arXiv:2002.06956}.

\bibitem[\protect\citename{Ouimet, }2020d]{arXiv:2001.08512}
Ouimet, F. 2020d.
\newblock A precise local limit theorem for the multinomial distribution and
  some applications.
\newblock {\em Preprint},  1--20.
\newblock \href{https://arxiv.org/abs/2001.08512}{arXiv:2001.08512}.

\bibitem[\protect\citename{Petrone, }1999a]{MR1703623}
Petrone, S. 1999a.
\newblock Bayesian density estimation using {B}ernstein polynomials.
\newblock {\em Canad. J. Statist.}, {\bf 27}(1), 105--126.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR1703623}{MR1703623}.

\bibitem[\protect\citename{Petrone, }1999b]{MR1712051}
Petrone, S. 1999b.
\newblock Random {B}ernstein polynomials.
\newblock {\em Scand. J. Statist.}, {\bf 26}(3), 373--393.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR1712051}{MR1712051}.

\bibitem[\protect\citename{Petrone \& Wasserman, }2002]{MR1881846}
Petrone, S., \& Wasserman, L. 2002.
\newblock Consistency of {B}ernstein polynomial posteriors.
\newblock {\em J. Roy. Statist. Soc. Ser. B}, {\bf 64}(1), 79--100.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR1881846}{MR1881846}.

\bibitem[\protect\citename{Stadtm\"uller, }1983]{MR0726014}
Stadtm\"uller, U. 1983.
\newblock Asymptotic distributions of smoothed histograms.
\newblock {\em Metrika}, {\bf 30}(3), 145--158.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR0726014}{MR0726014}.

\bibitem[\protect\citename{Stadtm\"uller, }1986]{MR0858109}
Stadtm\"uller, U. 1986.
\newblock Asymptotic properties of nonparametric curve estimates.
\newblock {\em Period. Math. Hungar.}, {\bf 17}(2), 83--108.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR0858109}{MR0858109}.

\bibitem[\protect\citename{Tenbusch, }1994]{MR1293514}
Tenbusch, A. 1994.
\newblock Two-dimensional {B}ernstein polynomial density estimators.
\newblock {\em Metrika}, {\bf 41}(3-4), 233--253.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR1293514}{MR1293514}.

\bibitem[\protect\citename{Tenbusch, }1997]{MR1437794}
Tenbusch, A. 1997.
\newblock Nonparametric curve estimation with {B}ernstein estimates.
\newblock {\em Metrika}, {\bf 45}(1), 1--30.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR1437794}{MR1437794}.

\bibitem[\protect\citename{Turnbull \& Ghosh, }2014]{MR3139345}
Turnbull, B.~C., \& Ghosh, S.~K. 2014.
\newblock Unimodal density estimation using {B}ernstein polynomials.
\newblock {\em Comput. Statist. Data Anal.}, {\bf 72}, 13--29.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR3139345}{MR3139345}.

\bibitem[\protect\citename{Vitale, }1975]{MR0397977}
Vitale, R.~A. 1975.
\newblock Bernstein polynomial approach to density function estimation.
\newblock {\em Pages  87--99 of:} {\em Statistical {I}nference and {R}elated
  {T}opics}.
\newblock Academic Press, New York.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR0397977}{MR0397977}.

\end{thebibliography}

\documentclass[final,3p,10pt,authoryear]{elsarticle}

\makeatletter
\def\ps@pprintTitle{%
 \let\@oddhead\@empty
 \let\@evenhead\@empty
 \def\@oddfoot{}%
 \let\@evenfoot\@oddfoot}
\makeatother

\usepackage{amssymb,amsthm,mathtools}
\usepackage[colorlinks]{hyperref}
\usepackage{natbib}

\usepackage{xcolor,dsfont,geometry}
\newtheorem{theorem}{Theorem}

\makeatletter \@addtoreset{equation}{section} \makeatother

\newcommand{\N}{\mathbb{N}}

\newcommand{\EE}{\mathbb{E}}

\newcommand{\bb}[1]{\boldsymbol{#1}}

\newcommand{\leqdef}{\vcentcolon=}

\newcommand{\blue}{\color{blue}}
\newcommand{\ind}{\mathds{1}}

\allowdisplaybreaks

\begin{document}

\begin{frontmatter}

    \title{Explicit formulas for the joint third and fourth central moments\\of the multinomial distribution}%

    \author[a1]{Fr\'ed\'eric Ouimet\texorpdfstring{\corref{cor1}\fnref{fn1}}{)}}%

    \address[a1]{California Institute of Technology, Pasadena, USA.}%

    \cortext[cor1]{Corresponding author}%
    \ead{ouimetfr@caltech.edu}%

    \begin{abstract}
        We give the first explicit formulas for the joint third and fourth central moments of the multinomial distribution, by differentiating the moment generating function. A general formula for the joint factorial moments was previously given in \cite{MR143299}.
    \end{abstract}

    \begin{keyword}
        multinomial distribution \sep simplex \sep central moments \sep third moment \sep fourth moment
        \MSC[2010]{Primary : 62E15 Secondary : 60E05}
    \end{keyword}

\end{frontmatter}

\vspace{-2mm}
\section{The multinomial distribution}\label{sec:intro}

    For any $d\in \N$, the $d$-dimensional (unit) simplex is defined by $\mathcal{S} \leqdef \big\{\bb{x}\in [0,1]^d : \sum_{i=1}^d x_i \leq 1\big\}$, and the probability mass function $\bb{k}\mapsto P_{\bb{k},m}(\bb{x})$ for $\bb{\xi} \leqdef (\xi_1,\xi_2,\dots,\xi_d) \sim \mathrm{Multinomial}\hspace{0.2mm}(m,\bb{x})$ is defined by
    \begin{equation}\label{eq:multinomial.probability}
        P_{\bb{k},m}(\bb{x}) \leqdef \frac{m!}{(m - \sum_{i=1}^d k_i)! \prod_{i=1}^d k_i!} \cdot (1 - \sum_{i=1}^d x_i)^{m - \sum_{i=1}^d k_i} \prod_{i=1}^d x_i^{k_i}, \quad \bb{k}\in \N_0^d \cap m\mathcal{S},
    \end{equation}
    where $m\in \N$ and $\bb{x}\in \mathcal{S}$.
    In this paper, our goal is to compute, for all $i,j,\ell,p\in \{1,2,\dots,d\}$,
    \begin{equation}
        \EE\big[(\xi_i - \EE[\xi_i])(\xi_j - \EE[\xi_j])(\xi_{\ell} - \EE[\xi_{\ell}])\big] \quad \text{and} \quad \EE\big[(\xi_i - \EE[\xi_i])(\xi_j - \EE[\xi_j])(\xi_{\ell} - \EE[\xi_{\ell}])(\xi_p - \EE[\xi_p])\big].
    \end{equation}

\section{Motivation}\label{sec:motivation}

    To the best of our knowledge, explicit formulas for the joint third and fourth central moments of the multinomial distribution have never been derived in the literature.
    These central moments can arise naturally, for example, when studying asymptotic properties, via Taylor expansions, of statistical estimators involving the multinomial distribution.
    For a given sequence of i.i.d.\ observations $\bb{X}_1,\bb{X}_2,\dots,\bb{X}_n$, two examples of such estimators are the Bernstein estimator for the cumulative distribution function
    \begin{equation}\label{eq:cdf.Bernstein.estimator}
        F_{n,m}^{\star}(\bb{x}) \leqdef \sum_{\bb{k}\in \N_0^d \cap m \mathcal{S}} \frac{1}{n} \sum_{i=1}^n \ind_{(-\bb{\infty},\frac{\bb{k}}{m}]}(\bb{X}_i) P_{\bb{k},m}(\bb{x}), \quad \bb{x}\in \mathcal{S}, ~m,n \in \N,
    \end{equation}
    and the Bernstein estimator for the density function (also called smoothed histogram)
    \begin{equation}\label{eq:histogram.estimator}
        \hat{f}_{n,m}(\bb{x}) \leqdef \sum_{\bb{k}\in \N_0^d \cap (m-1) \mathcal{S}} \frac{m^d}{n} \sum_{i=1}^n \ind_{(\frac{\bb{k}}{m}, \frac{\bb{k} + 1}{m}]}(\bb{X}_i) P_{\bb{k},m-1}(\bb{x}), \quad \bb{x}\in \mathcal{S}, ~m,n \in \N,
    \end{equation}
    over the $d$-dimensional simplex.
    Some of their asymptotic properties were investigated by \cite{MR0397977}, \cite{MR0638651}, \cite{MR0726014}, \cite{MR0791719}, \cite{MR0858109}, \cite{MR1437794}, \cite{MR1703623}, \cite{MR1712051}, \cite{MR1873330}, \cite{MR1881846}, \cite{MR1910059}, \cite{MR2068610}, \cite{MR2351744}, \cite{MR2395599}, \cite{MR2488150}, \cite{MR2662607}, \cite{MR2782409}, \cite{MR2960952}, \cite{MR2925964}, \cite{MR3174309}, \cite{MR3139345}, \cite{MR3412755}, \cite{MR3488598}, \cite{MR3740722}, \cite{MR3630225}, \cite{MR3983257} when $d = 1$, by \cite{MR1293514} when $d = 2$, and by \cite{arXiv:2002.07758,arXiv:2006.11756} for all $d\geq 1$, using a local limit theorem from \cite{arXiv:2001.08512} for the multinomial distribution (see also \cite{MR0478288}).
    The estimator \eqref{eq:histogram.estimator} is a discrete analogue of the Dirichlet kernel estimator introduced by \cite{doi:10.2307/2347365} and studied theoretically in \cite{MR1685301}, \cite{MR1718494}, \cite{MR1742101}, \cite{MR1985506} when $d = 1$ (among others), and in \cite{arXiv:2002.06956} for all $d\geq 1$.

\newpage
\section{Results}\label{sec:main.results}

    First, we compute the non-central moments of the multinomial distribution in \eqref{eq:multinomial.probability}.
    These moments were also calculated independently on a case-by-case basis in \cite{Newcomer2008phd,Newcomer_et_al_2008_tech_report}.

    \begin{theorem}[{\blue Non-central moments}]\label{thm:non.central.moments}
        Let $m^{(k)} \leqdef m (m-1) \dots (m-k+1)$ and let $\bb{\xi} \sim \mathrm{Multinomial}\hspace{0.3mm}(m,\bb{x})$.
        Then, for all $i,j,\ell,p\in \{1,2,\dots,d\}$,
        \begin{align}
            \EE[\xi_i]
            &= m \, x_i, \label{eq:lem:non.central.moments.eq.1} \\[2mm]
            \EE[\xi_i \xi_j]
            &= m^{(2)} x_i x_j + \ind_{\{i = j\}} m \, x_i, \label{eq:lem:non.central.moments.eq.2} \\[2mm]
            \EE[\xi_i \xi_j \xi_{\ell}]
            &= m^{(3)} x_i x_j x_{\ell} + \ind_{\{i = j = \ell\}} \, [3 \, m^{(2)} x_i^2 + m \, x_i] \label{eq:lem:non.central.moments.eq.3} \\[-0.5mm]
            &\quad+ \ind_{\{i = j \neq \ell \neq i\}} m^{(2)} x_i x_{\ell} + \ind_{\{i \neq j = \ell \neq i\}} m^{(2)} x_i x_j + \ind_{\{i \neq j \neq \ell = i\}} m^{(2)} x_j x_{\ell}, \notag \\[2mm]
            \EE[\xi_i \xi_j \xi_{\ell} \xi_p]
            &= m^{(4)} x_i x_j x_{\ell} x_p \label{eq:lem:non.central.moments.eq.4} \\[-1mm]
            &\quad+ \ind_{\{i = j \neq \ell \neq i\} \cap \{i \neq p \neq \ell\}} \, m^{(3)} x_i x_{\ell} x_p \notag \\
            &\quad+ \ind_{\{i \neq j = \ell \neq i\} \cap \{i \neq p \neq j\}} \, m^{(3)} x_i x_j x_p \notag \\
            &\quad+ \ind_{\{i \neq j \neq \ell = i\} \cap \{j \neq p \neq \ell\}} \, m^{(3)} x_j x_{\ell} x_p \notag \\[1mm]
            &\quad+ \ind_{\{i = j \neq \ell \neq i\} \cap \{i = p \neq \ell\}} \, [3 \, m^{(3)} x_i^2 x_{\ell} + m^{(2)} x_i x_{\ell}] \notag \\
            &\quad+ \ind_{\{i \neq j = \ell \neq i\} \cap \{i \neq p = j\}} \, [3 \, m^{(3)} x_i x_j^2 + m^{(2)} x_i x_j] \notag \\
            &\quad+ \ind_{\{i \neq j \neq \ell = i\} \cap \{j \neq p = \ell\}} \, [3 \, m^{(3)} x_j x_{\ell}^2 + m^{(2)} x_j x_{\ell}] \notag \\[1mm]
            &\quad+ \ind_{\{p = i\} \cap \{j \neq \ell \neq p \neq j\}} \, m^{(3)} x_i x_j x_{\ell} \notag \\
            &\quad+ \ind_{\{p = j\} \cap \{i \neq \ell \neq p \neq i\}} \, m^{(3)} x_i x_j x_{\ell} \notag \\
            &\quad+ \ind_{\{p = \ell\} \cap \{i \neq j \neq p \neq i\}} \, m^{(3)} x_i x_j x_{\ell} \notag \\[1mm]
            &\quad+ \ind_{\{i = j \neq \ell \neq i\} \cap \{i \neq p = \ell\}} \, [m^{(3)} x_i^2 x_{\ell} + m^{(3)} x_i x_{\ell}^2 + m^{(2)} x_i x_{\ell}] \notag \\
            &\quad+ \ind_{\{i \neq j = \ell \neq i\} \cap \{i = p \neq j\}} \, [m^{(3)} x_i x_j^2 + m^{(3)} x_i^2 x_j + m^{(2)} x_i x_j] \notag \\
            &\quad+ \ind_{\{i \neq j \neq \ell = i\} \cap \{j = p \neq \ell\}} \, [m^{(3)} x_j x_{\ell}^2 + m^{(3)} x_j^2 x_{\ell} + m^{(2)} x_j x_{\ell}] \notag \\[1mm]
            &\quad+ \ind_{\{i = j = \ell \neq p\}} \, [3 \, m^{(3)} x_i^2 x_p + m^{(2)} x_i x_p] \notag \\[1mm]
            &\quad+ \ind_{\{i = j = \ell = p\}} \, [6 \, m^{(3)} x_i^3 + 7 m^{(2)} x_i^2 + m x_i]. \notag
        \end{align}
    \end{theorem}

    \begin{proof}
        The moment generating function of $\bb{\xi}$ is
        \begin{equation}
            M_m(\bb{t}) = \bigg(1 - \sum_{i=1}^d x_i + \sum_{i=1}^d x_i \, e^{t_i}\bigg)^m.
        \end{equation}

        \vspace{-3mm}
        \noindent
        We have
        \begin{align}
            \frac{\partial}{\partial t_i} M_m(\bb{t})
            &= m M_{m-1}(\bb{t}) \, x_i \, e^{t_i}, \\[1mm]
            \frac{\partial^2}{\partial t_i \partial t_j} M_m(\bb{t})
            &= m^{(2)} M_{m-2}(\bb{t}) \, x_i \, e^{t_i} \, x_j \, e^{t_j} + \ind_{\{i = j\}} m M_{m-1}(\bb{t}) \, x_i \, e^{t_i}, \\[1mm]
            \frac{\partial^3}{\partial t_i \partial t_j \partial t_{\ell}} M_m(\bb{t})
            &= m^{(3)} M_{m-3}(\bb{t}) \, x_i \, e^{t_i} \, x_j \, e^{t_j} \, x_{\ell} \, e^{t_{\ell}} \\[-3mm]
            &\quad+ \ind_{\{i = j \neq \ell \neq i\}} m^{(2)} M_{m-2}(\bb{t}) \, x_i \, e^{t_i} \, x_{\ell} \, e^{t_{\ell}} \notag \\
            &\quad+ \ind_{\{i \neq j = \ell \neq i\}} m^{(2)} M_{m-2}(\bb{t}) \, x_i \, e^{t_i} \, x_j \, e^{t_j} \notag \\
            &\quad+ \ind_{\{i \neq j \neq \ell = i\}} m^{(2)} M_{m-2}(\bb{t}) \, x_j \, e^{t_j} \, x_{\ell} \, e^{t_{\ell}} \notag \\
            &\quad+ \ind_{\{i = j = \ell\}} \, [3 \, m^{(2)} M_{m-2}(\bb{t}) \, x_i^2 \, e^{2 t_i} + m M_{m-1}(\bb{t}) \, x_i \, e^{t_i}], \notag \\[1mm]
            \frac{\partial^4}{\partial t_i \partial t_j \partial t_{\ell} \partial t_p} M_m(\bb{t})
            &= m^{(4)} M_{m-4}(\bb{t}) \, x_i \, e^{t_i} \, x_j \, e^{t_j} \, x_{\ell} \, e^{t_{\ell}} \, x_p \, e^{t_p} \\[-2mm]
            &\quad+ \ind_{\{i = j \neq \ell \neq i\} \cap \{i \neq p \neq \ell\}} m^{(3)} M_{m-3}(\bb{t}) \, x_i \, e^{t_i} \, x_{\ell} \, e^{t_{\ell}} \, x_p \, e^{t_p} \notag \\
            &\quad+ \ind_{\{i \neq j = \ell \neq i\} \cap \{i \neq p \neq j\}} m^{(3)} M_{m-3}(\bb{t}) \, x_i \, e^{t_i} \, x_j e^{t_j} \, x_p \, e^{t_p} \notag \\
            &\quad+ \ind_{\{i \neq j \neq \ell = i\} \cap \{j \neq p \neq \ell\}} m^{(3)} M_{m-3}(\bb{t}) \, x_j \, e^{t_j} \, x_{\ell} \, e^{t_{\ell}} \, x_p \, e^{t_p} \notag \\[1mm]
            &\quad+ \ind_{\{i = j \neq \ell \neq i\} \cap \{i = p \neq \ell\}} \, \left[\hspace{-1mm}
                \begin{array}{l}
                    3 \, m^{(3)} M_{m-3}(\bb{t}) \, x_i^2 \, e^{2 t_i} \, x_{\ell} \, e^{t_{\ell}} \\
                    + m^{(2)} M_{m-2}(\bb{t}) \, x_i \, e^{t_i} \, x_{\ell} \, e^{t_{\ell}}
                \end{array}
                \hspace{-1mm}\right] \notag \\
            &\quad+ \ind_{\{i \neq j = \ell \neq i\} \cap \{i \neq p = j\}} \, \left[\hspace{-1mm}
                \begin{array}{l}
                    3 \, m^{(3)} M_{m-3}(\bb{t}) \, x_i \, e^{t_i} \, x_j^2 \, e^{2 t_j} \\
                    + m^{(2)} M_{m-2}(\bb{t}) \, x_i e^{t_i} \, x_j \, e^{t_j}
                \end{array}
                \hspace{-1mm}\right] \notag \\
            &\quad+ \ind_{\{i \neq j \neq \ell = i\} \cap \{j \neq p = \ell\}} \, \left[\hspace{-1mm}
                \begin{array}{l}
                    3 \, m^{(3)} M_{m-3}(\bb{t}) x_j \, e^{t_j} \, x_{\ell}^2 \, e^{2 t_{\ell}} \\
                    + m^{(2)} M_{m-2}(\bb{t}) \, x_j \, e^{t_j} \, x_{\ell} \, e^{t_{\ell}}
                \end{array}
                \hspace{-1mm}\right] \notag \\[1mm]
            &\quad+ \ind_{\{p = i\} \cap \{j \neq \ell \neq p \neq j\}} m^{(3)} M_{m-3}(\bb{t}) \, x_i \, e^{t_i} \, x_j \, e^{t_j} \, x_{\ell} \, e^{t_{\ell}} \notag \\
            &\quad+ \ind_{\{p = j\} \cap \{i \neq \ell \neq p \neq i\}} m^{(3)} M_{m-3}(\bb{t}) \, x_i \, e^{t_i} \, x_j \, e^{t_j} \, x_{\ell} \, e^{t_{\ell}} \notag \\
            &\quad+ \ind_{\{p = \ell\} \cap \{i \neq j \neq p \neq i\}} m^{(3)} M_{m-3}(\bb{t}) \, x_i \, e^{t_i} \, x_j \, e^{t_j} \, x_{\ell} \, e^{t_{\ell}} \notag \\[1mm]
            &\quad+ \ind_{\{i = j \neq \ell \neq i\} \cap \{i \neq p = \ell\}} \, \left[\hspace{-1mm}
                \begin{array}{l}
                    m^{(3)} M_{m-3}(\bb{t}) \, x_i^2 \, e^{2 t_i} \, x_{\ell} \, e^{t_{\ell}} \\
                    + m^{(3)} M_{m-3}(\bb{t}) \, x_i \, e^{t_i} \, x_{\ell}^2 \, e^{2 t_{\ell}} \\
                    + m^{(2)} M_{m-2}(\bb{t}) \, x_i \, e^{t_i} \, x_{\ell} \, e^{t_{\ell}}
                \end{array}
                \hspace{-1.5mm}\right] \notag \\
            &\quad+ \ind_{\{i \neq j = \ell \neq i\} \cap \{i = p \neq j\}} \, \left[\hspace{-1mm}
                \begin{array}{l}
                    m^{(3)} M_{m-3}(\bb{t}) \, x_i \, e^{t_i} \, x_j^2 \, e^{2 t_j} \\
                    + m^{(3)} M_{m-3}(\bb{t}) \, x_i^2 \, e^{2 t_i} \, x_j \, e^{t_j} \\
                    + m^{(2)} M_{m-2}(\bb{t}) \, x_i \, e^{t_i} \, x_j \, e^{t_j}
                \end{array}
                \hspace{-1.5mm}\right] \notag \\
            &\quad+ \ind_{\{i \neq j \neq \ell = i\} \cap \{j = p \neq \ell\}} \, \left[\hspace{-1mm}
                \begin{array}{l}
                    m^{(3)} M_{m-3}(\bb{t}) \, x_j \, e^{t_j} \, x_{\ell}^2 \, e^{2 t_{\ell}} \\
                    + m^{(3)} M_{m-3}(\bb{t}) \, x_j^2 \, e^{2 t_j} \, x_{\ell} \, e^{t_{\ell}} \\
                    + m^{(2)} M_{m-2}(\bb{t}) \, x_j \, e^{t_j} \, x_{\ell} \, e^{t_{\ell}}
                \end{array}
                \hspace{-1.5mm}\right] \notag \\[1mm]
            &\quad+ \ind_{\{i = j = \ell \neq p\}} \, \left[\hspace{-1mm}
                \begin{array}{l}
                    3 \, m^{(3)} M_{m-3}(\bb{t}) \, x_i^2 \, e^{2 t_i} \, x_p \, e^{t_p} \\
                    + m^{(2)} M_{m-2}(\bb{t}) \, x_i \, e^{t_i} \, x_p \, e^{t_p}
                \end{array}
                \hspace{-1mm}\right] \notag \\[1mm]
            &\quad+ \ind_{\{i = j = \ell = p\}} \, \left[\hspace{-1mm}
                \begin{array}{l}
                    6 m^{(3)} M_{m-3}(\bb{t}) \, x_i^3 \, e^{3 t_i} \\
                    + 7 m^{(2)} M_{m-2}(\bb{t}) \, x_i^2 \, e^{2 t_i} \\
                    + m M_{m-1}(\bb{t}) \, x_i \, e^{t_i}
                \end{array}
                \hspace{-1.5mm}\right]. \notag
        \end{align}
        By taking $\bb{t} = \bb{0}$, we get the conclusion.
    \end{proof}

    With some algebraic manipulations and a careful analysis, we can now obtain the central moments.

    \begin{theorem}[{\blue Central moments}]\label{thm:central.moments}
        Let $\bb{\xi} \sim \mathrm{Multinomial}\hspace{0.3mm}(m,\bb{x})$, then, for all $i,j,\ell,p\in \{1,2,\dots,d\}$,
        \begin{align}
            &\EE\big[\xi_i - \EE[\xi_i]\big] = 0, \label{eq:thm:central.moments.eq.1} \\[2mm]
            &\EE\big[(\xi_i - \EE[\xi_i])(\xi_j - \EE[\xi_j])\big] = m \, (x_i \ind_{\{i = j\}} - x_i x_j), \label{eq:thm:central.moments.eq.2} \\[2mm]
            &\EE\big[(\xi_i - \EE[\xi_i])(\xi_j - \EE[\xi_j])(\xi_{\ell} - \EE[\xi_{\ell}])\big] \notag \\[1mm]
            &\hspace{5mm}= m \, \big(2 x_i x_j x_{\ell} - \ind_{\{i = j\}} x_i x_{\ell} - \ind_{\{j = \ell\}} x_i x_j - \ind_{\{i = \ell\}} x_j x_{\ell} + \ind_{\{i = j = \ell\}} x_i\big), \label{eq:thm:central.moments.eq.3} \\[2mm]
            &\EE\big[(\xi_i - \EE[\xi_i])(\xi_j - \EE[\xi_j])(\xi_{\ell} - \EE[\xi_{\ell}])(\xi_p - \EE[\xi_p])\big] \notag \\[1mm]
            &\hspace{5mm}= (3 m^2 - 6 m) \, x_i x_j x_{\ell} x_p - (12 m^2 - 12 m) \, \ind_{\{i = j = \ell = p\}} x_i^3 \label{eq:thm:central.moments.eq.4} \\[0.5mm]
            &\hspace{5mm}\quad+ m^2 \left\{\hspace{-1mm}
                \begin{array}{l}
                    \ind_{\{i = j\}} x_i x_{\ell} x_p + \ind_{\{i = \ell\}} x_j x_{\ell} x_p + \ind_{\{i = p\}} x_i x_j x_{\ell} \\
                    \ind_{\{j = \ell\}} x_i x_j x_p + \ind_{\{j = p\}} x_i x_j x_{\ell} + \ind_{\{\ell = p\}} x_i x_j x_{\ell}
                \end{array}
                \hspace{-1mm}\right\} \notag \\[0.5mm]
            &\hspace{5mm}\quad+ (3 m^2 - 7 m) \, \ind_{\{i = j = \ell = p\}} x_i^2 + m \, \ind_{\{i = j = \ell = p\}} x_i \notag \\[0.5mm]
            &\hspace{5mm}\quad- (6 m^2 - 6 m) \left\{\hspace{-1mm}
                \begin{array}{l}
                    \ind_{\{i = j = \ell \neq p\}} x_i^2 x_p + \ind_{\{i = j = p \neq \ell\}} x_i^2 x_{\ell} \\
                    + \ind_{\{i = \ell = p \neq j\}} x_j x_{\ell}^2 + \ind_{\{j = \ell = p \neq i\}} x_i x_j^2
                \end{array}
                \hspace{-1mm}\right\} \notag \\[0.5mm]
            &\hspace{5mm}\quad- (2 m^2 - 2 m) \left\{\hspace{-1mm}
                \begin{array}{l}
                    \ind_{\{\ell \neq i = j \neq p\}} x_i x_{\ell} x_p + \ind_{\{j \neq i = \ell \neq p\}} x_j x_{\ell} x_p + \ind_{\{j \neq i = p \neq \ell\}} x_i x_j x_{\ell} \\
                    \ind_{\{i \neq j = \ell \neq p\}} x_i x_j x_p + \ind_{\{i \neq j = p \neq \ell\}} x_i x_j x_{\ell} + \ind_{\{i \neq \ell = p \neq j\}} x_i x_j x_{\ell}
                \end{array}
                \hspace{-1mm}\right\} \notag \\[0.5mm]
            &\hspace{5mm}\quad+ (m^2 - m) \, \big\{\ind_{\{i = j \neq \ell = p\}} x_i x_{\ell} + \ind_{\{i = p \neq j = \ell\}} x_i x_j + \ind_{\{i = \ell \neq j = p\}} x_j x_{\ell}\big\} \notag \\[0.5mm]
            &\hspace{5mm}\quad- m \, \big\{\ind_{\{i = j = \ell \neq p\}} x_i x_p + \ind_{\{i = j = p \neq \ell\}} x_i x_{\ell} + \ind_{\{i = \ell = p \neq j\}} x_j x_{\ell} + \ind_{\{j = \ell = p \neq i\}} x_i x_j\big\}. \notag
        \end{align}
    \end{theorem}

    \begin{proof}
        The expression \eqref{eq:thm:central.moments.eq.2} for the covariance follows directly from \eqref{eq:lem:non.central.moments.eq.1} and \eqref{eq:lem:non.central.moments.eq.2}.
        For the third central moments, \eqref{eq:lem:non.central.moments.eq.1}, \eqref{eq:lem:non.central.moments.eq.2} and \eqref{eq:lem:non.central.moments.eq.3} yield
        \begin{align*}
            &\EE\big[(\xi_i - \EE[\xi_i])(\xi_j - \EE[\xi_j])(\xi_{\ell} - \EE[\xi_{\ell}])\big] \\[1mm]
            &\hspace{5mm}= \EE[\xi_i \xi_j \xi_{\ell}] - m \, \EE[\xi_i \xi_j] \, x_{\ell} - m \, \EE[\xi_i \xi_{\ell}] \, x_j - m \, \EE[\xi_j \xi_{\ell}] \, x_i \\[0.5mm]
            &\hspace{5mm}\quad+ m^2 \EE[\xi_i] \, x_j x_{\ell} + m^2 \EE[\xi_j] \, x_i x_{\ell} + m^2 \EE[\xi_{\ell}] \, x_i x_j - m^3 x_i x_j x_{\ell} \\[0.5mm]
            &\hspace{5mm}= \left\{\hspace{-1mm}
                \begin{array}{l}
                    m^{(3)} \, x_i x_j x_{\ell} + \ind_{\{i = j = \ell\}} \, [3 \, m^{(2)} x_i^2 + m \, x_i] \\[0.5mm]
                    + \ind_{\{i = j \neq \ell \neq i\}} m^{(2)} x_i x_{\ell} + \ind_{\{i \neq j = \ell \neq i\}} m^{(2)} x_i x_j + \ind_{\{i \neq j \neq \ell = i\}} m^{(2)} x_j x_{\ell}
                \end{array}
                \hspace{-1mm}\right\} \\[0.5mm]
            &\hspace{5mm}\quad- m \, \big\{m^{(2)} x_i x_j + \ind_{\{i = j\}} m \, x_i\big\} x_{\ell} - m \, \big\{m^{(2)} x_i x_{\ell} + \ind_{\{i = {\ell}\}} m \, x_i\big\} \, x_j \\[0.5mm]
            &\hspace{5mm}\quad- m \, \big\{m^{(2)} x_j x_{\ell} + \ind_{\{j = \ell\}} m \, x_j\big\} x_i + 2 m^3 x_i x_j x_{\ell} \\[0.5mm]
            &\hspace{5mm}= 2 m \, x_i x_j x_{\ell} - m \, \ind_{\{i = j\}} x_i x_{\ell} - m \, \ind_{\{j = \ell\}} x_i x_j - m \, \ind_{\{i = \ell\}} x_j x_{\ell} + m \, \ind_{\{i = j = \ell\}} x_i \\[0.5mm]
            &\hspace{5mm}= m \, \big(2 x_i x_j x_{\ell} - \ind_{\{i = j\}} x_i x_{\ell} - \ind_{\{j = \ell\}} x_i x_j - \ind_{\{i = \ell\}} x_j x_{\ell} + \ind_{\{i = j = \ell\}} x_i\big).
        \end{align*}
        Similarly, by \eqref{eq:lem:non.central.moments.eq.1}, \eqref{eq:lem:non.central.moments.eq.2}, \eqref{eq:lem:non.central.moments.eq.3} and \eqref{eq:lem:non.central.moments.eq.4}, we have
        \begin{align*}
            &\EE\big[(\xi_i - \EE[\xi_i])(\xi_j - \EE[\xi_j])(\xi_{\ell} - \EE[\xi_{\ell}])(\xi_p - \EE[\xi_p])\big] \\[1mm]
            &\hspace{5mm}= \EE[\xi_i \xi_j \xi_{\ell} \xi_p] - m \, \EE[\xi_i \xi_j \xi_{\ell}] \, x_p - m \, \EE[\xi_i \xi_j \xi_p] \, x_{\ell} - m \, \EE[\xi_i \xi_{\ell} \xi_p] \, x_j - m \, \EE[\xi_j \xi_{\ell} \xi_p] \, x_i \\[0.75mm]
            &\hspace{5mm}\quad+ m^2 \EE[\xi_i \xi_j] \, x_{\ell} x_p + m^2 \EE[\xi_i \xi_{\ell}] \, x_j x_p + m^2 \EE[\xi_i \xi_p] \, x_j x_{\ell} \\[0.75mm]
            &\hspace{5mm}\quad+ m^2 \EE[\xi_j \xi_{\ell}] \, x_i x_p + m^2 \EE[\xi_j \xi_p] \, x_i x_{\ell} + m^2 \EE[\xi_{\ell} \xi_p] \, x_i x_j \\[0.75mm]
            &\hspace{5mm}\quad- m^3 \EE[\xi_i] \, x_i x_j x_{\ell} - m^3 \EE[\xi_j] \, x_i x_{\ell} x_p - m^3 \EE[\xi_{\ell}] \, x_i x_j x_p - m^3 \EE[\xi_p] \, x_i x_j x_{\ell} + m^4 x_i x_j x_{\ell} x_p \\[1mm]
            &\hspace{5mm}= \left\{\hspace{-1mm}
                \begin{array}{l}
                    m^{(4)} x_i x_j x_{\ell} x_p \notag \\
                    + \ind_{\{i = j \neq \ell \neq i\} \cap \{i \neq p \neq \ell\}} \, m^{(3)} x_i x_{\ell} x_p \notag \\
                    + \ind_{\{i \neq j = \ell \neq i\} \cap \{i \neq p \neq j\}} \, m^{(3)} x_i x_j x_p \notag \\
                    + \ind_{\{i \neq j \neq \ell = i\} \cap \{j \neq p \neq \ell\}} \, m^{(3)} x_j x_{\ell} x_p \notag \\[1.5mm]
                    + \ind_{\{i = j \neq \ell \neq i\} \cap \{i = p \neq \ell\}} \, [3 \, m^{(3)} x_i^2 x_{\ell} + m^{(2)} x_i x_{\ell}] \notag \\
                    + \ind_{\{i \neq j = \ell \neq i\} \cap \{i \neq p = j\}} \, [3 \, m^{(3)} x_i x_j^2 + m^{(2)} x_i x_j] \notag \\
                    + \ind_{\{i \neq j \neq \ell = i\} \cap \{j \neq p = \ell\}} \, [3 \, m^{(3)} x_j x_{\ell}^2 + m^{(2)} x_j x_{\ell}] \notag \\[1.5mm]
                    + \ind_{\{p = i\} \cap \{j \neq \ell \neq p \neq j\}} \, m^{(3)} x_i x_j x_{\ell} \notag \\
                    + \ind_{\{p = j\} \cap \{i \neq \ell \neq p \neq i\}} \, m^{(3)} x_i x_j x_{\ell} \notag \\
                    + \ind_{\{p = \ell\} \cap \{i \neq j \neq p \neq i\}} \, m^{(3)} x_i x_j x_{\ell} \notag \\[1.5mm]
                    + \ind_{\{i = j \neq \ell \neq i\} \cap \{i \neq p = \ell\}} \, [m^{(3)} x_i^2 x_{\ell} + m^{(3)} x_i x_{\ell}^2 + m^{(2)} x_i x_{\ell}] \notag \\
                    + \ind_{\{i \neq j = \ell \neq i\} \cap \{i = p \neq j\}} \, [m^{(3)} x_i x_j^2 + m^{(3)} x_i^2 x_j + m^{(2)} x_i x_j] \notag \\
                    + \ind_{\{i \neq j \neq \ell = i\} \cap \{j = p \neq \ell\}} \, [m^{(3)} x_j x_{\ell}^2 + m^{(3)} x_j^2 x_{\ell} + m^{(2)} x_j x_{\ell}] \notag \\[1.5mm]
                    + \ind_{\{i = j = \ell \neq p\}} \, [3 \, m^{(3)} x_i^2 x_p + m^{(2)} x_i x_p] \notag \\[1.5mm]
                    + \ind_{\{i = j = \ell = p\}} \, [6 m^{(3)} x_i^3 + 7 m^{(2)} x_i^2 + m x_i]
                \end{array}
                \hspace{-1mm}\right\} \\[1mm]
            &\hspace{5mm}\quad- m \left\{\hspace{-1mm}
                \begin{array}{l}
                    m^{(3)} \, x_i x_j x_{\ell} + \ind_{\{i = j = \ell\}} \, [3 \, m^{(2)} x_i^2 + m \, x_i] \\[0.5mm]
                    + \ind_{\{i = j \neq \ell \neq i\}} m^{(2)} x_i x_{\ell} + \ind_{\{i \neq j = \ell \neq i\}} m^{(2)} x_i x_j + \ind_{\{i \neq j \neq \ell = i\}} m^{(2)} x_j x_{\ell}
                \end{array}
                \hspace{-1mm}\right\} x_p \\[1mm]
            &\hspace{5mm}\quad- m \left\{\hspace{-1mm}
                \begin{array}{l}
                    m^{(3)} \, x_i x_j x_p + \ind_{\{i = j = p\}} \, [3 \, m^{(2)} x_i^2 + m \, x_i] \\[0.5mm]
                    + \ind_{\{i = j \neq p \neq i\}} m^{(2)} x_i x_p + \ind_{\{i \neq j = p \neq i\}} m^{(2)} x_i x_j + \ind_{\{i \neq j \neq p = i\}} m^{(2)} x_j x_p
                \end{array}
                \hspace{-1mm}\right\} x_{\ell} \\[1mm]
            &\hspace{5mm}\quad- m \left\{\hspace{-1mm}
                \begin{array}{l}
                    m^{(3)} \, x_i x_{\ell} x_p + \ind_{\{i = \ell = p\}} \, [3 \, m^{(2)} x_i^2 + m \, x_i] \\[0.5mm]
                    + \ind_{\{i = \ell \neq p \neq i\}} m^{(2)} x_i x_p + \ind_{\{i \neq \ell = p \neq i\}} m^{(2)} x_i x_{\ell} + \ind_{\{i \neq \ell \neq p = i\}} m^{(2)} x_{\ell} x_p
                \end{array}
                \hspace{-1mm}\right\} x_j \\[1mm]
            &\hspace{5mm}\quad- m \left\{\hspace{-1mm}
                \begin{array}{l}
                    m^{(3)} \, x_j x_{\ell} x_p + \ind_{\{j = \ell = p\}} \, [3 \, m^{(2)} x_j^2 + m \, x_j] \\[0.5mm]
                    + \ind_{\{j = \ell \neq p \neq j\}} m^{(2)} x_j x_p + \ind_{\{j \neq \ell = p \neq j\}} m^{(2)} x_j x_{\ell} + \ind_{\{j \neq \ell \neq p = j\}} m^{(2)} x_{\ell} x_p
                \end{array}
                \hspace{-1mm}\right\} x_i \\[1mm]
            &\hspace{5mm}\quad+ m^2 \big\{m^{(2)} x_i x_j + \ind_{\{i = j\}} m \, x_i\big\}  x_{\ell} x_p + m^2 \big\{m^{(2)} x_i x_{\ell} + \ind_{\{i = \ell\}} m \, x_i\big\} x_j x_p \\[1mm]
            &\hspace{5mm}\quad+ m^2 \big\{m^{(2)} x_i x_p + \ind_{\{i = p\}} m \, x_i\big\} x_j x_{\ell} + m^2 \big\{m^{(2)} x_j x_{\ell} + \ind_{\{j = \ell\}} m \, x_j\big\} x_i x_p \\[1mm]
            &\hspace{5mm}\quad+ m^2 \big\{m^{(2)} x_j x_p + \ind_{\{j = p\}} m \, x_j\big\} x_i x_{\ell} + m^2 \big\{m^{(2)} x_{\ell} x_p + \ind_{\{\ell = p\}} m \, x_{\ell}\big\} x_i x_j - 3 m^4 x_i x_j x_{\ell} x_p.
        \end{align*}
        We notice that all the terms with powers $m^3$ and $m^4$ cancel out with each other.
        The above is thus
        \begin{align*}
            &= \left\{\hspace{-1mm}
                \begin{array}{l}
                    (11 m^2 - 6 m) \, x_i x_j x_{\ell} x_p \notag \\[1mm]
                    + (- 3 m^2 + 2 m) \, \ind_{\{i = j \neq \ell \neq i\} \cap \{i \neq p \neq \ell\}} x_i x_{\ell} x_p \notag \\
                    + (- 3 m^2 + 2 m) \, \ind_{\{i \neq j = \ell \neq i\} \cap \{i \neq p \neq j\}} x_i x_j x_p \notag \\
                    + (- 3 m^2 + 2 m) \, \ind_{\{i \neq j \neq \ell = i\} \cap \{j \neq p \neq \ell\}} x_j x_{\ell} x_p \notag \\[1.5mm]
                    + (- 9 m^2 + 6 m) \, \ind_{\{i = j \neq \ell \neq i\} \cap \{i = p \neq \ell\}} \, x_i^2 x_{\ell} \notag \\
                    + (- 9 m^2 + 6 m) \, \ind_{\{i \neq j = \ell \neq i\} \cap \{i \neq p = j\}} \, x_i x_j^2 \notag \\
                    + (- 9 m^2 + 6 m) \, \ind_{\{i \neq j \neq \ell = i\} \cap \{j \neq p = \ell\}} \, x_j x_{\ell}^2 \notag \\[1.5mm]
                    + (1 m^2 - 1 m) \, \ind_{\{i = j \neq \ell \neq i\} \cap \{i = p \neq \ell\}} \, x_i x_{\ell} \notag \\
                    + (1 m^2 - 1 m) \, \ind_{\{i \neq j = \ell \neq i\} \cap \{i \neq p = j\}} \, x_i x_j \notag \\
                    + (1 m^2 - 1 m) \, \ind_{\{i \neq j \neq \ell = i\} \cap \{j \neq p = \ell\}} \, x_j x_{\ell} \notag \\[1.5mm]
                    + (- 3 m^2 + 2 m) \, \ind_{\{p = i\} \cap \{j \neq \ell \neq p \neq j\}} x_i x_j x_{\ell} \notag \\
                    + (- 3 m^2 + 2 m) \, \ind_{\{p = j\} \cap \{i \neq \ell \neq p \neq i\}} x_i x_j x_{\ell} \notag \\
                    + (- 3 m^2 + 2 m) \, \ind_{\{p = \ell\} \cap \{i \neq j \neq p \neq i\}} x_i x_j x_{\ell} \notag \\[1.5mm]
                    + (- 3 m^2 + 2 m) \, \ind_{\{i = j \neq \ell \neq i\} \cap \{i \neq p = \ell\}} \, x_i^2 x_{\ell} \notag \\
                    + (- 3 m^2 + 2 m) \, \ind_{\{i \neq j = \ell \neq i\} \cap \{i = p \neq j\}} \, x_i x_j^2 \notag \\
                    + (- 3 m^2 + 2 m) \, \ind_{\{i \neq j \neq \ell = i\} \cap \{j = p \neq \ell\}} \, x_j x_{\ell}^2 \notag \\[1.5mm]
                    + (- 3 m^2 + 2 m) \, \ind_{\{i = j \neq \ell \neq i\} \cap \{i \neq p = \ell\}} \, x_i x_{\ell}^2 \notag \\
                    + (- 3 m^2 + 2 m) \, \ind_{\{i \neq j = \ell \neq i\} \cap \{i = p \neq j\}} \, x_i^2 x_j \notag \\
                    + (- 3 m^2 + 2 m) \, \ind_{\{i \neq j \neq \ell = i\} \cap \{j = p \neq \ell\}} \, x_j^2 x_{\ell} \notag \\[1.5mm]
                    + (1 m^2 - 1 m) \, \ind_{\{i = j \neq \ell \neq i\} \cap \{i \neq p = \ell\}} \, x_i x_{\ell} \notag \\
                    + (1 m^2 - 1 m) \, \ind_{\{i \neq j = \ell \neq i\} \cap \{i = p \neq j\}} \, x_i x_j \notag \\
                    + (1 m^2 - 1 m) \, \ind_{\{i \neq j \neq \ell = i\} \cap \{j = p \neq \ell\}} \, x_j x_{\ell} \notag \\[1.5mm]
                    + (- 9 m^2 + 6 m) \, \ind_{\{i = j = \ell \neq p\}} \, x_i^2 x_p \notag \\[1.5mm]
                    + (1 m^2 - 1 m) \, \ind_{\{i = j = \ell \neq p\}} \, x_i x_p \notag \\[1.5mm]
                    + (- 18 m^2 + 12 m) \, \ind_{\{i = j = \ell = p\}} \, x_i^3 \notag \\[1.5mm]
                    + (7 m^2 - 7 m) \, \ind_{\{i = j = \ell = p\}} \, x_i^2 \notag \\[1mm]
                    + 1 m \, \ind_{\{i = j = \ell = p\}} x_i
                \end{array}
                \hspace{-1mm}\right\} \\[1mm]
            &\quad+ \left\{\hspace{-1mm}
                \begin{array}{l}
                    - 2 m^2 \, x_i x_j x_{\ell} x_p + 3 m^2 \, \ind_{\{i = j = \ell\}} \, x_i^2 x_p - 1 m^2 \, \ind_{\{i = j = \ell\}} \, x_i x_p \\[0.5mm]
                    + 1 m^2 \, \ind_{\{i = j \neq \ell \neq i\}} x_i x_{\ell} x_p + 1 m^2 \, \ind_{\{i \neq j = \ell \neq i\}} x_i x_j x_p + 1 m^2 \, \ind_{\{i \neq j \neq \ell = i\}} x_j x_{\ell} x_p
                \end{array}
                \hspace{-1mm}\right\} \\[1mm]
            &\quad+ \left\{\hspace{-1mm}
                \begin{array}{l}
                    - 2 m^2 \, x_i x_j x_{\ell} x_p + 3 m^2 \, \ind_{\{i = j = p\}} \, x_i^2 x_{\ell} - 1 m^2 \, \ind_{\{i = j = p\}} \, x_i x_{\ell} \\[0.5mm]
                    + 1 m^2 \, \ind_{\{i = j \neq p \neq i\}} x_i x_{\ell} x_p + 1 m^2 \, \ind_{\{i \neq j = p \neq i\}} x_i x_j x_{\ell} + 1 m^2 \, \ind_{\{i \neq j \neq p = i\}} x_j x_{\ell} x_p
                \end{array}
                \hspace{-1mm}\right\} \\[1mm]
            &\quad+ \left\{\hspace{-1mm}
                \begin{array}{l}
                    - 2 m^2 \, x_i x_j x_{\ell} x_p + 3 m^2 \, \ind_{\{i = \ell = p\}} \, x_i^2 x_j - 1 m^2 \, \ind_{\{i = \ell = p\}} \, x_i x_j \\[0.5mm]
                    + 1 m^2 \, \ind_{\{i = \ell \neq p \neq i\}} x_i x_j x_p + 1 m^2 \, \ind_{\{i \neq \ell = p \neq i\}} x_i x_j x_{\ell} + 1 m^2 \, \ind_{\{i \neq \ell \neq p = i\}} x_j x_{\ell} x_p
                \end{array}
                \hspace{-1mm}\right\} \\[1mm]
            &\quad+ \left\{\hspace{-1mm}
                \begin{array}{l}
                    - 2 m^2 \, x_i x_j x_{\ell} x_p + 3 m^2 \, \ind_{\{j = \ell = p\}} \, x_i x_j^2 - 1 m^2 \, \ind_{\{j = \ell = p\}} \, x_i x_j \\[0.5mm]
                    + 1 m^2 \, \ind_{\{j = \ell \neq p \neq j\}} x_i x_j x_p + 1 m^2 \, \ind_{\{j \neq \ell = p \neq j\}} x_i x_j x_{\ell} + 1 m^2 \, \ind_{\{j \neq \ell \neq p = j\}} x_i x_{\ell} x_p
                \end{array}
                \hspace{-1mm}\right\} \\[2mm]
            &= (3 m^2 - 6 m) \, x_i x_j x_{\ell} x_p - (12 m^2 - 12 m) \, \ind_{\{i = j = \ell = p\}} x_i^3  \\
            &\quad+ m^2 \left\{\hspace{-1mm}
                \begin{array}{l}
                    \ind_{\{i = j\}} x_i x_{\ell} x_p + \ind_{\{i = \ell\}} x_j x_{\ell} x_p + \ind_{\{i = p\}} x_i x_j x_{\ell} \\
                    \ind_{\{j = \ell\}} x_i x_j x_p + \ind_{\{j = p\}} x_i x_j x_{\ell} + \ind_{\{\ell = p\}} x_i x_j x_{\ell}
                \end{array}
                \hspace{-1mm}\right\} \\[0.5mm]
            &\quad+ (3 m^2 - 7 m) \, \ind_{\{i = j = \ell = p\}} x_i^2 + m \, \ind_{\{i = j = \ell = p\}} x_i \\[0.5mm]
            &\quad- (6 m^2 - 6 m) \left\{\hspace{-1mm}
                \begin{array}{l}
                    \ind_{\{i = j = \ell \neq p\}} x_i^2 x_p + \ind_{\{i = j = p \neq \ell\}} x_i^2 x_{\ell} \\
                    + \ind_{\{i = \ell = p \neq j\}} x_j x_{\ell}^2 + \ind_{\{j = \ell = p \neq i\}} x_i x_j^2
                \end{array}
                \hspace{-1mm}\right\} \notag \\[0.5mm]
            &\quad- (2 m^2 - 2 m) \left\{\hspace{-1mm}
                \begin{array}{l}
                    \ind_{\{\ell \neq i = j \neq p\}} x_i x_{\ell} x_p + \ind_{\{j \neq i = \ell \neq p\}} x_j x_{\ell} x_p + \ind_{\{j \neq i = p \neq \ell\}} x_i x_j x_{\ell} \\
                    \ind_{\{i \neq j = \ell \neq p\}} x_i x_j x_p + \ind_{\{i \neq j = p \neq \ell\}} x_i x_j x_{\ell} + \ind_{\{i \neq \ell = p \neq j\}} x_i x_j x_{\ell}
                \end{array}
                \hspace{-1mm}\right\} \\[0.5mm]
            &\quad+ (m^2 - m) \, \big\{\ind_{\{i = j \neq \ell = p\}} x_i x_{\ell} + \ind_{\{i = p \neq j = \ell\}} x_i x_j + \ind_{\{i = \ell \neq j = p\}} x_j x_{\ell}\big\} \\[0.5mm]
            &\quad- m \, \big\{\ind_{\{i = j = \ell \neq p\}} x_i x_p + \ind_{\{i = j = p \neq \ell\}} x_i x_{\ell} + \ind_{\{i = \ell = p \neq j\}} x_j x_{\ell} + \ind_{\{j = \ell = p \neq i\}} x_i x_j\big\}.
        \end{align*}
        This ends the proof.
    \end{proof}

\section{Conclusion}

    In this short paper, we found explicit expressions for the third and fourth central and non-central moments of the multinomial distribution.
    The joint factorial moments were previously calculated in the literature (by \cite{MR143299}), but not the third and fourth central moments.

\section*{Abbreviations}

    i.i.d.\ : independent and identically distributed

\section*{Declarations}

    \subsection{Availability of data and material}

        Not applicable.

    \subsection{Funding}

        The author is supported by a postdoctoral fellowship from the NSERC (PDF) and a supplement from the FRQNT (B3X).

    \subsection{Competing interests}

        The author declares no conflict of interest.

    \subsection{Author's contributions}

        All contributions were made by the sole author of the article, Fr\'ed\'eric Ouimet.

    \subsection{Acknowledgments}

        Not applicable.

%
%

\nocite{MR3825458}
\nocite{Ouimet2019phd}

\bibliographystyle{authordate1}
\bibliography{Ouimet_2020_third_fourth_moments_multinomial_bib}

\end{document}